Maturity (years) Price (per $100 face value) | 1 | 2 | 3 | 4 | 5 94.52 89.68 85.40 81.65 78.35 The yield to maturity for theThe following table summarizes prices of various default-free zero-coupon bonds (expressed as a percentage of face value):
Answer #1

YTM of 4 year zero coupon = (FV/PV)^(1/n) - 1

FV = FACE VALUE = 100, PRICE = PV = 81.65

YTM of 4 year zero coupon =(100/81.65)^(1/4) - 1 = 5.20%

Answer : e : 5.20% (Thumbs up please)